# MSTAR: Box-free Multi-query Scene Text Retrieval with Attention Recycling

### v1

@@ Abstract
Scene text retrieval has made significant progress with the assistance of accurate text localization. However, existing approaches typically require costly bounding box annotations for training. Besides, they mostly adopt a customized retrieval strategy but struggle to unify various types of queries to meet diverse retrieval needs. To address these issues, we introduce Muti-query Scene Text retrieval with Attention Recycling (MSTAR), a box-free approach for scene text retrieval. It incorporates progressive vision embedding to dynamically capture the multi-grained representation of texts and harmonizes free-style text queries with style-aware instructions. Additionally, a multi-instance matching module is integrated to enhance vision-language alignment. Furthermore, we build the Multi-Query Text Retrieval (MQTR) dataset, the first benchmark designed to evaluate the multi-query scene text retrieval capability of models, comprising four query types and $16k$ images. Extensive experiments demonstrate the superiority of our method across seven public datasets and the MQTR dataset. Notably, MSTAR marginally surpasses the previous state-of-the-art model by 6.4% in MAP on Total-Text while eliminating box annotation costs. Moreover, on the MQTR benchmark, MSTAR significantly outperforms the previous models by an average of 8.5%. The code and datasets are available at https://github.com/yingift/MSTAR.
@@ 1 Introduction
Figure 1: (a) MSTAR achieves scene text retrieval without the aid of box annotations. (b) Image-text matching experiments with VLM  [19]. Detailed text instances like “welcome to beautiful” and “old florida” in the image receive lower matching scores. While manually covering salient text regions which receive the higher scores, the model can adaptively recognize the detailed text.
@@
Scene text appears almost everywhere in daily life and is an essential ingredient for image-text searching [6]. Traditional scene text retrieval [9] typically aims to search images based on word or phrase query. However, the real-life retrieval needs could be diverse. For instance, people often search an article with several key words, which cannot be fulfilled by a single word. Additionally, the non-ocr visual semantics is also vital for scene text searching. In this work, we study the multi-query scene text retrieval that aims to handle queries of various types within a unified model. This could facilitate applications like searching disambiguation [31, 16] and visual document indexing [45, 4].
@@
Off-the-shelf scene text retrieval methods [9, 37, 38] achieve text retrieval by explicitly localizing the text and matching the query with scene text instances, as shown in Fig. 1 (a). A straightforward solution is the text spotting methods [22, 20, 46]. They first spot the text in images and match it using edit distance. However, the two divided stages could cause error accumulation between text spotting and query matching. To mitigate this, detection-based methods [37, 40, 38] represent detected ROIs in dense embedding, integrating text detection and similarity learning into an end-to-end framework. Recently, FDP [48] leverages bounding boxes to guide CLIP [30] in focusing on text regions to achieve accurate retrieval. Despite these advancements, they require expensive box annotations for training. As different retrieval tasks require varying labels, it is very costly to obtain multiple types of bounding boxes, i.e., word-level, text-line, and common object bounding boxes.
@@
Recently, the large-scale box-free pretraining of Vision-Language Models (VLMs) [18, 49] has shown impressive capability on various tasks [1, 50, 47]. To apply box-free methods for scene text retrieval, we conduct image-text matching experiments, as shown in Fig.  1 (b). The observations reveal that box-free models tend to overlook detailed text instances. Moreover, manually masking salient text regions mitigates this issue by enabling the model to adaptively capture image details. More observations are supplemented in the appendix. Inspired by these, we introduce a box-free model termed Muti-query Scene Text retrieval with Attention Recycling (MSTAR). It starts with leveraging pretrained VLMs for scene text retrieval. To better capture detailed scene text features, the progressive vision embedding is designed to shift attention from salient regions to insalient areas with less attention iteratively. To support diverse retrieval queries, text queries are encoded by the multi-modal encoder with style-aware instructions. A multi-instance matching module is then designed to establish the cross-modal alignment. In this way, MSTAR seamlessly unifies diverse text queries and aligns image-query embeddings without the need of any positional supervision.
@@
To evaluate the performance of multi-query retrieval, we have carefully built a Multi-Query Text Retrieval (MQTR) benchmark. Beyond traditional word and phrase queries, this dataset introduces two additional, valuable retrieval settings: combined query and semantic query. The combined query comprises several discontinuous key texts (words or phrases) to enable more precise retrieval. Semantic queries, on the other hand, are image descriptions that require understanding both scene text and its non-ocr visual context. In total, the MQTR dataset includes four styles of queries (word, phrase, combined, and semantic) and $16000$ images.
@@
Experiments reveal that existing models struggle to simultaneously handle four types of query on the challenging MQTR dataset. In contrast, our proposed MSTAR performs well in multi-query retrieval. Notably, MSTAR impressively surpasses previous works by an average of 8.5% in MAP. Additionally, on seven public retrieval datasets, MSTAR demonstrates competitive performance compared to state-of-the-art box-based methods while significantly reducing annotation costs. Specifically, MSTAR outperforms FDP [48] by 6.4% in MAP on the widely used Total-Text dataset.
@@ 2 Related Work
Existing scene text retrieval datasets primarily focus on single-type retrieval such as word and phrase. The IIIT Scene Text Retrieval dataset [27] is a large-scale dataset dedicated to word retrieval comprising 10k images. The COCOText Retrieval dataset is derived from 7k natural images in the COCOText dataset [36]. In addition, several smaller but well-annotated datasets [39, 7] are commonly utilized for evaluating word retrieval performance. The Chinese Street View Text Retrieval dataset [37] contains 23 Chinese queries and 1,667 scene text images from Chinese street views. Besides these word retrieval datasets, the Phrase-level Scene Text Retrieval [48] dataset includes 36 frequently used phrase queries and 1,080 images. The CSVTRv2 [38] dataset supports partial and text-line queries. However, these datasets do not comprehensively support the evaluation of tasks such as combined retrieval and semantic retrieval. In constrast, our MQTR dataset can support four types of query to satisfy diverse needs in real-world applications.
@@
Existing methods generally adopt a paradigm that first localize text regions and then match with the query. In the early attempts, Mishra et al.[27] proposed identifying approximate character locations and indexing words using spatial constraints. More recent approaches try to integrate the two process into an end-to-end trainable framework. Gomez et al.[9] proposed to combine the detected proposals with the Pyramidal Histogram of Characters [2]. Wang et al.[37] unified the text detector with a cross-modal similarity model into an end-to-end framework. Its updated version [38] proposed the RankMIL and DPMA algorithms to address the partial scene text retrieval problem. Wen et al.[40] transformed cross-modal similarity into uni-modal similarity using image templates. Zeng et al.[48] utilized CLIP [30] for scene text retrieval and incorporated box supervision to localize text regions. However, these methods require expensive bounding box annotations for training.
In addition to the specifically designed retrieval methods, text spotters can also be applied to retrieval tasks. They first spot text instances from images and then rank with edit distance. Traditional text spotters with boundary supervision [22, 20, 11] can achieve accurate results. There are also point-supervised [29, 23] and transcription-only supervised methods [35, 41, 42], which yield limited performance. Above all, the separation of text spotting and query matching often leads to error accumulation, and text spotters struggle to handle multi-query retrieval.
Unlike above methods, our MSTAR achieves text retrieval without the bounding-box supervision and harmonize various data labels for multi-query scene text retrieval.
@@ 3.1 Overview
The overall architecture of MSTAR is depicted in Fig.  2. Given a scene text image, the vision encoder extracts image features $f_{\text{0}}$, and the Progressive Vision Embedding module progressively captures the scene text features as vision embeddings $E_{\text{V}}$. Simultaneously, text queries are encoded as text embeddings $E_{\text{T}}$ by the multi-modal encoder from BLIP2 [19] with style-aware instructions. The $E_{\text{V}}$ and $E_{\text{T}}$ are then fed into the Multi-Instance Matching module to align the cross-modal embeddings optimized with contrastive loss. Additionally, a reranking process is incorporated by inputting both image features and text queries into the multi-modal encoder to compute one-to-one matching scores. During training, MSTAR is optimized using both contrastive loss and image-text matching loss. For inference, image-text pairs are initially ranked by the cosine similarity, and the top K images are further matched with reranking process.
@@ 3.2 Progressive Vision Embedding
Vision-Language Models (VLMs) pretrained on image-caption pairs often focus more on salient visual concepts such as a red circle [33, 44]. However, the detailed and subtle scene text instances typically appear in various insalient regions of natural scenes. As depicted in Fig.  1 (b), the “old florida” and “welcome to beautiful” are overlooked as the model focus on more salient regions. This problem leads to a high miss rate for small text scenarios in scene text retrieval tasks. To mitigate this, we propose a Progressive Vision Embedding (PVE) approach to extract visual embeddings.
@@
Since the model tends to focus on salient image features (visualized in Fig. 2), $E_{\text{V}}^{0}$ struggles to fully capture the detailed text instance representation. To force the model to focus on less salient features, we propose the Salient Attention Shift (SAS) module. Motivated by observations in Sec. 1, the SAS uses a mask-attention layer to automatically shift image attention. Unlike previous methods  [5] that use ground truth as supervision, the mask in our approach is derived from the cross-attention layers of the multi-modal encoder $\psi$. Concretely, we first calculate the mean of the cross-attention weights of $\psi$ as the cross-attention map $C_{\text{t-1}}$. Then $C_{\text{t-1}}$ is binarized with a binarization algorithm $\sigma$ and inverted to obtain a binary mask $M_{\text{t-1}}$. This is formulated in Eq. 1.
@@ 3.3 Instruction Aware Text Representation
where $\psi$ denotes the multi-modal encoder, $T_{\text{i}}$ represents instructions and $T_{\text{Q}}$ denotes text queries. The instructions prompt the $\psi$ to distinguish query types. As shown in Fig. 2, queries of each style are encoded into different representation space during training. To speed up training, all of the text queries (words, phrases, combined, and semantic queries) paired to the image are encoded altogether. The $\psi$ encodes the queries into text embeddings $E_{\text{T}}\in\mathbb{R}^{\text{N}\times\text{d}}$ which consist of single-word embeddings $E_{\text{w}}\in\mathbb{R}^{N_{\text{w}}\times\text{d}}$ and multi-word embeddings $E_{\text{m}}\in\mathbb{R}^{N_{\text{m}}\times\text{d}}$. The N is the total number of text queries paired to the image. The N is the total number of text queries paired to the image. The $N_{\text{w}}$ and $N_{\text{m}}$ denotes the number of single-word queries and multi-word queries, respectively.
3.4 Muti-Instance Matching
After obtaining the vision embeddings $E_{\text{V}}\in\mathbb{R}^{\text{(T+1)Q}\times\text{d}}$ and text embeddings $E_{\text{T}}\in\mathbb{R}^{\text{N}\times\text{d}}$, the problem is to build the one-to-one alignment for the multi-type and multi-instance embeddings. The previous study either adopts aggregate the multiple vision embeddings into one embedding [19] or adopts the late interaction [15, 8]. However, due to the implicit matching mechanism, these strategies need massive training for vision-language alignment.
[…]
In the multi-word branch, since the multi-word queries contain abundant semantic information, a light-weight cross attention layer is used to aggregate the vision features under text constraint in the second branch. This process is formulated as Eq. 4.
@@ 3.5 Optimization
where $\alpha$ is a hyper-parameter to maintain numerical approximation equivalence of the two losses. Since the number of queries are usually greater than the number of images, $\alpha$ is set to 1.5 in our implementation. The image-text matching process simultaneously feeds image features and text queries into the multi-modal encoder $\psi$. An image-text matching score is then computed using a linear layer with a two-cell output. This score is optimized using a cross-entropy matching loss $\mathcal{L}_{\text{m}}$. The overall loss is the sum of the $\mathcal{L}_{\text{c}}$ and $\mathcal{L}_{\text{m}}$.
@@ 4.1 Multi-Query Text Retrieval Dataset
To comprehensively evaluate the performance of models on multi-query scene text retrieval, we carefully build the Multi-Query Text Retrieval (MQTR) dataset. The MQTR dataset includes four sub-tasks: word, phrase, combined, and semantic retrieval. The construction of this dataset leverages well-annotated public datasets [36, 21, 13, 14, 7, 34, 24, 48], along with images obtained from Google Image Search. The word, phrase, and combined subsets each contain 5,000 images and the 200 most frequently occurring queries. The semantic subset consists of 1,000 images and 25 queries collected from the web. The semantic subset was manually collected with 10-15 positive images and an equal number of hard negative samples for each query. The hard negatives samples refer to images that contain three types of objects: (1) visual elements with semantics similar to the query (e.g., an apple and the word “apple”), (2) text instances with similar shapes, and (3) text instances with similar meanings. The inclusion of hard negatives poses additional challenges by introducing visually and textually confounding samples, thus assessing the capacity of retrieval models to distinguish visual semantics from OCR-based semantics. As demonstrated in Tab.  1, our MQTR dataset is the first comprehensive benchmark to support four types of query in scene text retrieval. You can refer to the appendix for more construction details and images samples from the datasets.

### v2

@@ Abstract
Scene text retrieval has made significant progress with the assistance of accurate text localization. However, existing approaches typically require costly bounding box annotations for training. Besides, they mostly adopt a customized retrieval strategy but struggle to unify various types of queries to meet diverse retrieval needs. To address these issues, we introduce Multi-query Scene Text retrieval with Attention Recycling (MSTAR), a box-free approach for scene text retrieval. It incorporates progressive vision embedding to dynamically capture the multi-grained representation of texts and harmonizes free-style text queries with style-aware instructions. Additionally, a multi-instance matching module is integrated to enhance vision-language alignment. Furthermore, we build the Multi-Query Text Retrieval (MQTR) dataset, the first benchmark designed to evaluate the multi-query scene text retrieval capability of models, comprising four query types and $16k$ images. Extensive experiments demonstrate the superiority of our method across seven public datasets and the MQTR dataset. Notably, MSTAR marginally surpasses the previous state-of-the-art model by 6.4% in MAP on Total-Text while eliminating box annotation costs. Moreover, on the MQTR benchmark, MSTAR significantly outperforms the previous models by an average of 8.5%. The code and datasets are available at https://github.com/yingift/MSTAR.
@@ 1 Introduction
Figure 1: (a) MSTAR achieves scene text retrieval without the aid of box annotations. (b) Image-text matching experiments with VLM [li2023blip]. Detailed text instances like “welcome to beautiful” and “old florida” in the image receive lower matching scores. While manually covering salient text regions which receive the higher scores, the model can adaptively recognize the detailed text.
@@
Scene text appears almost everywhere in daily life and is an essential ingredient for image-text searching [cheng2022vista]. Traditional scene text retrieval [gomez2018single] typically aims to search images based on word or phrase queries and could benefit applications such as handwritten signature retrieval [pavenet2025zhang, zhang2025capturing] and key frame extraction [song2019text]. However, real-life retrieval needs could be diverse. For instance, people often search for an article with several key words, which cannot be fulfilled with a single word. Additionally, non-ocr visual semantics is also vital for scene text searching. In this work, we study the multi-query scene text retrieval that aims to handle queries of various types within a unified model. This could facilitate applications such as searching disambiguation [schroeder2020structured, kritharoula2023large] and visual document indexing [yang2017smart, caffagni2024wiki].
@@
Off-the-shelf scene text retrieval methods [gomez2018single, wang2021scene, wang2024partial] achieve text retrieval by explicitly localizing the text and matching the query with scene text instances, as shown in Fig. 1 (a). A straightforward solution is the text spotting methods [liu2021abcnet, liao2020mask, ye2023deepsolo]. They first spot the text in images and match it using edit distance. However, the two divided stages could cause error accumulation between text spotting and query matching. To mitigate this, detection-based methods [wang2021scene, wen2023visual, wang2024partial] represent detected ROIs in dense embedding, integrating text detection and similarity learning into an end-to-end framework. Recently, FDP [zeng2024focus] leverages bounding boxes to guide CLIP [radford2021learning] in focusing on text regions to achieve accurate retrieval. Despite these advancements, they require expensive box annotations for training. As different retrieval tasks require varying labels, it is very costly to obtain multiple types of bounding boxes, i.e., word-level, text-line, and common object bounding boxes.
@@
Recently, the large-scale box-free pre-training of Vision-Language Models (VLMs) [li2022blip, zhai2023sigmoid] has shown impressive capability on various tasks [aberdam2023clipter, zhao2023clip4str, yu2023turning]. To apply box-free methods for scene text retrieval, we conduct image-text matching experiments, as shown in Fig.  1 (b). The observations reveal that box-free models tend to overlook detailed text instances. Moreover, manually masking salient text regions mitigates this issue by enabling the model to adaptively capture image details. More observations are supplemented in the appendix. Inspired by these, we introduce a box-free model termed Multi-query Scene Text retrieval with Attention Recycling (MSTAR). It starts with leveraging pre-trained VLMs for scene text retrieval. To better capture detailed scene text features, the progressive vision embedding is designed to shift attention from salient regions to insalient areas with less attention iteratively. To support diverse retrieval queries, text queries are encoded by the multi-modal encoder with style-aware instructions. A multi-instance matching module is then designed to establish the cross-modal alignment. In this way, MSTAR seamlessly unifies diverse text queries and aligns image-query embeddings without the need of any positional supervision.
@@
To evaluate the performance of multi-query retrieval, we have carefully built a Multi-Query Text Retrieval (MQTR) benchmark. Beyond traditional word and phrase queries, this dataset introduces two additional, valuable retrieval settings: combined query and semantic query. The combined query comprises several discontinuous key texts (words or phrases) to enable more precise retrieval. Semantic queries, on the other hand, are image descriptions that require understanding both scene text and its non-ocr visual context [xie2024pdfwukong]. In total, the MQTR dataset includes four styles of queries (word, phrase, combined, and semantic) and $16000$ images.
@@
Experiments reveal that existing models struggle to simultaneously handle four types of query on the challenging MQTR dataset. In contrast, our proposed MSTAR performs well in multi-query retrieval. Notably, MSTAR impressively surpasses previous work by an average of 8.5% in MAP. Additionally, on seven public retrieval datasets, MSTAR demonstrates competitive performance compared to state-of-the-art box-based methods while significantly reducing annotation costs. Specifically, MSTAR outperforms FDP [zeng2024focus] by 6.4% in MAP on the widely used Total-Text dataset.
@@ 2 Related Work
Existing scene text retrieval datasets primarily focus on single-type retrieval such as word and phrase. The IIIT Scene Text Retrieval dataset [mishra2013image] is a large-scale dataset dedicated to word retrieval comprising 10k images. The COCOText Retrieval dataset is derived from 7k natural images in the COCOText dataset [veit2016coco]. In addition, several smaller but well-annotated datasets [wang2011end, ch2020total] are commonly utilized to evaluate word retrieval performance. The Chinese Street View Text Retrieval dataset [wang2021scene] contains 23 Chinese queries and 1,667 scene text images from Chinese street views. Besides these word retrieval datasets, the Phrase-level Scene Text Retrieval [zeng2024focus] dataset includes 36 frequently used phrase queries and 1,080 images. The CSVTRv2 [wang2024partial] dataset supports partial and text-line queries. However, these datasets do not comprehensively support the evaluation of tasks such as combined retrieval and semantic retrieval. In contrast, our MQTR dataset can support four types of query to satisfy diverse needs in real-world applications.
@@
Existing methods generally adopt a paradigm that first localizes text regions and then matches with the query. In the early attempts, Mishra et al.[mishra2013image] proposed identifying approximate character locations and indexing words using spatial constraints. More recent approaches try to integrate the two processes into an end-to-end trainable framework. Gomez et al.[gomez2018single] proposed to combine the detected proposals with the Pyramidal Histogram of Characters [almazan2014word]. Wang et al.[wang2021scene] unified the text detector with a cross-modal similarity model into an end-to-end framework. Its updated version [wang2024partial] proposed the RankMIL and DPMA algorithms to address the partial scene text retrieval problem. Wen et al.[wen2023visual] transformed cross-modal similarity into uni-modal similarity using image templates. Zeng et al.[zeng2024focus] utilized CLIP [radford2021learning] for scene text retrieval and incorporated box supervision to localize text regions. However, these methods require expensive bounding box annotations for training.
In addition to the specifically designed retrieval methods, text spotters can also be applied to retrieval tasks. They first spot text instances from images and then rank with edit distance. Traditional text spotters with boundary supervision [liu2021abcnet, liao2020mask, xie2022toward] can achieve accurate results. There are also point-supervised methods [peng2022spts, liu2023spts] and transcription-only supervised methods [tang2022you, wen2024twist, wu2025wecromcl], which yield limited performance. Above all, the separation of text spotting and query matching often leads to error accumulation, and text spotters struggle to handle multi-query retrieval.
Unlike above methods, our MSTAR achieves text retrieval without the bounding-box supervision and harmonizes various data labels for multi-query scene text retrieval.
@@ 3.1 Overview
The overall architecture of MSTAR is depicted in Fig.  2. Given a scene text image, the vision encoder extracts image features $f_{\text{0}}$, and the Progressive Vision Embedding module progressively captures the scene text features as vision embeddings $E_{\text{V}}$. Simultaneously, text queries are encoded as text embeddings $E_{\text{T}}$ by the multi-modal encoder from BLIP-2 [li2023blip] with style-aware instructions. The $E_{\text{V}}$ and $E_{\text{T}}$ are then fed into the Multi-Instance Matching module to align the cross-modal embeddings optimized with contrastive loss. Additionally, a re-ranking process is incorporated by inputting both image features and text queries into the multi-modal encoder to compute one-to-one matching scores. During training, MSTAR is optimized using both contrastive loss and image-text matching loss. For inference, image-text pairs are initially ranked by the cosine similarity, and the top K images are further matched with re-ranking process.
@@ 3.2 Progressive Vision Embedding
Vision-Language Models (VLMs) pre-trained on image-caption pairs often focus more on salient visual concepts such as a red circle [shtedritski2023does, yang2024fine]. However, the detailed and subtle scene text instances typically appear in various insalient regions of natural scenes. As depicted in Fig.  1 (b), the “old florida” and “welcome to beautiful” are overlooked as the model focus on more salient regions. This problem leads to a high miss rate for small text scenarios in scene text retrieval tasks. To mitigate this, we propose a Progressive Vision Embedding (PVE) approach to extract visual embeddings.
@@
Since the model tends to focus on salient image features (visualized in Fig. 2), $E_{\text{V}}^{0}$ struggles to fully capture the detailed text instance representation. To force the model to focus on less salient features, we propose the Salient Attention Shift (SAS) module. Motivated by observations in Sec. 1, the SAS uses a mask-attention layer to automatically shift image attention. Unlike previous methods  [cheng2022masked] that use ground truth as supervision, the mask in our approach is derived from the cross-attention layers of the multi-modal encoder $\psi$. Concretely, we first calculate the mean of the cross-attention weights of $\psi$ as the cross-attention map $C_{\text{t-1}}$. Then $C_{\text{t-1}}$ is binarized with a binarization algorithm $\sigma$ and inverted to obtain a binary mask $M_{\text{t-1}}$. This is formulated in Eq. 1.
@@ 3.3 Instruction Aware Text Representation
where $\psi$ denotes the multi-modal encoder, $T_{\text{i}}$ represents instructions and $T_{\text{Q}}$ denotes text queries. The instructions prompt the $\psi$ to distinguish query types. As shown in Fig. 2, queries of each style are encoded into different representation space during training. To speed up training, all of the text queries (words, phrases, combined, and semantic queries) paired to the image are encoded altogether. The $\psi$ encodes the queries into text embeddings $E_{\text{T}}\in\mathbb{R}^{\text{N}\times\text{d}}$ which consist of single-word embeddings $E_{\text{w}}\in\mathbb{R}^{N_{\text{w}}\times\text{d}}$ and multi-word embeddings $E_{\text{m}}\in\mathbb{R}^{N_{\text{m}}\times\text{d}}$. The N is the total number of text queries paired to the image. The $N_{\text{w}}$ and $N_{\text{m}}$ denote the number of single-word queries and multi-word queries, respectively.
3.4 Multi-Instance Matching
After obtaining the vision embeddings $E_{\text{V}}\in\mathbb{R}^{\text{(T+1)Q}\times\text{d}}$ and text embeddings $E_{\text{T}}\in\mathbb{R}^{\text{N}\times\text{d}}$, the problem is to build the one-to-one alignment for the multi-type and multi-instance embeddings. The previous study either aggregate the multiple vision embeddings into one embedding [li2023blip] or adopts the late interaction [khattab2020colbert, faysse2024colpali]. However, due to the implicit matching mechanism, these strategies need massive training for vision-language alignment.
To mitigate these, we propose the Multi-Instance Matching (MIM) module to explicitly assign the one-to-one matching relations for vision-language embeddings. MIM comprises two parallel branches for processing single-word and multi-word queries, respectively. In the single-word branch, the Hungarian matching algorithm [kuhn1955hungarian] is exploited to explicitly assign the one-to-one matching relation between $E_{\text{w}}\in\mathbb{R}^{N_{\text{w}}\times\text{d}}$ and $E_{\text{V}}\in\mathbb{R}^{\text{(T+1)Q}\times\text{d}}$. Specifically, we first construct a cosine similarity matrix of size $N_{\text{w}}\times\text{(T+1)Q}$. Since $N_{\text{w}}$ is typically unequal to (T+1)Q, we pad the matrix with zeros to create a square matrix. Finally, the first $N_{\text{w}}$ rows of the results are used for one-to-one correspondences.
In the multi-word branch, since the multi-word queries contain abundant semantic information, a light-weight cross-attention layer is used to aggregate the vision features under text constraint in the second branch. This process is formulated as Eq. 4.
@@ 3.5 Optimization
where $\alpha$ is a hyperparameter to maintain the numerical approximation equivalence of the two losses. Since the number of queries is usually greater than the number of images, $\alpha$ is set to 1.5 in our implementation. The image-text matching process simultaneously feeds image features and text queries into the multi-modal encoder $\psi$. An image-text matching score is then computed using a linear layer with a two-cell output. This score is optimized using a cross-entropy matching loss $\mathcal{L}_{\text{m}}$. The overall loss is the sum of $\mathcal{L}_{\text{c}}$ and $\mathcal{L}_{\text{m}}$.
@@ 4.1 Multi-Query Text Retrieval Dataset
To comprehensively evaluate the performance of models on multi-query scene text retrieval, we carefully build the Multi-Query Text Retrieval (MQTR) dataset. The MQTR dataset includes four sub-tasks: word, phrase, combined, and semantic retrieval. The construction of this dataset leverages well-annotated public datasets [veit2016coco, liu2019curved, karatzas2013icdar, karatzas2015icdar, ch2020total, sidorov2020textcaps, long2023icdar, zeng2024focus], along with images obtained from Google Image Search. The word, phrase, and combined subsets each contain 5,000 images and the 200 most frequently occurring queries. The semantic subset consists of 1,000 images and 25 queries collected from the web. The semantic subset was manually collected with 10-15 positive images and an equal number of hard negative samples for each query. The hard negatives samples refer to images that contain three types of objects: (1) visual elements with semantics similar to the query (e.g., an apple and the word “apple”), (2) text instances with similar shapes, and (3) text instances with similar meanings. The inclusion of hard negatives poses additional challenges by introducing visually and textually confounding samples, thus assessing the capacity of retrieval models to distinguish visual semantics from OCR-based semantics. As demonstrated in Tab.  1, our MQTR dataset is the first comprehensive benchmark to support four types of query in scene text retrieval. You can refer to the appendix for more construction details and images samples from the datasets.